OpenStudy (he66666):

Is this set of vectors impossible to row reduce? [6 4 -2 4], [2 0 0 1], [3 2 -1 2], [5 6 -3 2], [0 4 -2 -1] I can't seem to row reduce it.. so I'm assuming it's not possible?

OpenStudy (he66666):

\[\left[\begin{matrix}2 & 1 &1&3&1 \\ 2 &0 &1&3&2\\ -2&0&-1&-3&-2\\2&1&1&-1&-3\end{matrix}\right]\] This is what I got to so far by first subtracting row1 by row4. And then I added row2 and row3. Finally, I added row3 and row4.

OpenStudy (amistre64):

row reducing is just using elementary row operations to reduce things down to 1s and 0s; or something cclose to it

OpenStudy (amistre64):

are those your column vectors?

OpenStudy (he66666):

The vectors in the questions are I think row vectors since it stated that it was R2(1x2 matrices)

OpenStudy (he66666):

Yes, I did use the elementary row operations but I ended up having the "first" leading 1 in the fourth row instead of the first row.

OpenStudy (amistre64):

you can move rows up and down as needed

OpenStudy (amistre64):

the system: 2x + 3y = 4 x - 2y = 2 has the same solutions as: x - 2y = 2 2x + 3y = 4

OpenStudy (amistre64):

im still not sure how your post gets setup tho

OpenStudy (he66666):

Oh, I forgot about how you can interchange the rows. Thanks a lot amistre!

OpenStudy (amistre64):

youre welcome